OpenStudy (anonymous):

please help George is trying to make some decisions about his business and has come up with P = (x - 100)(x + 20), where P represents his profit margin A)at he will make $100 if he sells no DVDS B) That the most amount of money he can make is $100. C) That in order to make zero profit he must sell 100 DVDs. D) That in order to make the most profit he must sell 100 DVDs

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

can you help me with a few more questions its a final exam and I don't know any of it

OpenStudy (anonymous):

Selena played 32 games of chess. If she won x games, which expression represents her losses? A) 32x B) 32 + x C) 32 - x D) 32 x

OpenStudy (anonymous):

total games 32 he won x so the lost is 32-x

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which expression represents how many years are in x months? A) 12x B) 12 + x C) x 12 D) 12 x

OpenStudy (anonymous):

12x

OpenStudy (anonymous):

number of months in x year = 12x

OpenStudy (anonymous):

A manufacturer uses a tank to supply water to his machines while they are working. When the machines are on the tank loses 1.5 gallons of water an hour at the same time 1 gallon of water is pumped into the tank from a well. Write an equation that models the amount of water (w) in the tank for any hour (h) that the machines are on. The tank starts full with 150 gallons of water in it. A) W = 1t + 150 B) W = -.5t + 150 C) W = 150t - 0.5 D) W = -1.5t + 150

OpenStudy (anonymous):

but number of year in x month= x/12

OpenStudy (anonymous):

w=150+1t-1.5t=150-.5t

OpenStudy (anonymous):

so the answer is B

OpenStudy (anonymous):

Alan, Bill, and Charlie are brothers. Bill is x years old, Alan is 1 year younger, and Charlie is 1 year older. . The expression that correctly shows the PRODUCT of the boys' ages is A) x3 B) 3x3 C) x3 + x D) x3 - x

OpenStudy (anonymous):

x(x-1)(x+1) (x^2-x)(x+1) x^3+x^2-x^2-x x^3-x

OpenStudy (anonymous):

so whats the answer d?

OpenStudy (anonymous):

yes